OpenStudy (anonymous):

Find the points on the curve y=x^3-2x+1 at which the tangent inclines with the negative direction of X-axis is at 135.

OpenStudy (anonymous):

"the tangent inclines with the negative direction of X-axis is at 135" means that it makes an angle of 135° or that it passes through (0,-135)?

OpenStudy (anonymous):

it makes an angle of 135°

OpenStudy (anonymous):

So, if it makes an angle of 135° with the negative direction of the x axis, it makes an agle of 180°-135°=45° with the positive direction, which means that the tangent has a slope of tan(45°) = 1

OpenStudy (turingtest):

I think you should draw the picture to get a better feel for this

OpenStudy (anonymous):

To find the point of the function where the tangent has a definite slope, you need to have the derivative of the funcion, which is y'(x) = 3x^2-2. So\[\large3x^2-2 = 1\rightarrow 3x^2 = 3 \rightarrow x^2 = 1 \rightarrow x = \pm 1\]

OpenStudy (anonymous):

thatas totally right,ty for your effort @Recursing